# Canonicalization for Unreproducible Builds in Java

### v1

@@ 7.3. Fixing Unreproducibility
Ren et al. (Ren et al., 2022, 2019, 2018) propose techniques to localize and fix unreproducibility in Debian packages. They localize the file in the Debian package that causes unreproducibility and then suggest a patch based on a history of patches for fixing unreproducible builds. This approach is specific for Debian packages and does not apply to Java artifacts. Moreover, their approach to detect unreproducibility is different from ours as mentioned in section 7.2.
@@
Xiong et al. (Xiong et al., 2022) classifies fixing unreproducibility into 3 categories - 1) remediation by patching source code, third-party dependency, or build script 2) controlling which means interception of non-deterministic build instructions and returning pre-defined values for them 3) interpretation which means documenting the reason for unreproducibility. The authors perform these fixes automatically using a tool called JavaBEPFix which is not available to the public. Hence, we cannot evaluate the effectiveness of their tool on our dataset.
@@
Dietrich et al. (Dietrich et al., 2024b) show that only a 25% (out of 14,156) of the binaries in Maven ecosystem are actually bit by bit identical. These binaries come from 4 different registries - Maven, Google Assured Open Source Software, RedHat, and Oracle build-from-source. They propose 4 levels of binary equivalence to compare binaries which becomes increasingly lenient as we go from level 1 to level 4. First is to compare binaries bit by bit and this is the strictest level of binary equivalence check. Second is to compare binaries and ignore differences that have no semantic effect (for example, @Deprecated). OSS-Rebuild fits into the second level of checking binary equivalence. Third is to convert binaries into a common format and then compare them. jNorm fits into the third level of checking binary equivalence Finally, fourth is to compare binaries using a similarity metric. In the paper, they use tlsh (Oliver et al., 2013) to compare hash distance between two binaries; lesser the hash distance, more similar the binaries are. This work is similar to ours as they also evaluate effectiveness of jNorm over thousands of pairs of JVM bytecode (Dietrich et al., 2024a) to fix unreproducibility issues. However, we consider more causes of unreproducibility other than ones in JVM bytecode.
Finally, strip-nondeterminism  [^38] https://salsa.debian.org/reproducible-builds/strip-nondeterminism is a tool the Debian project that strips out timestamps and file system order from archives. However, it only takes care of these 2 causes of unreproducibility and does not cover the other causes of unreproducibility that we have discussed in this paper.
@@ 8. Conclusion
In this paper, we have presented the first large-scale study of build unreproducibility in Java. We have identified the six main causes of unreproducibility with underlying root causes: 1) build manifest, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps. Leveraging the same dataset, we have evaluated the effectiveness of canonicalization as a mitigation strategy to fix unreproducibility. Our results have shown that 26.89% of unreproducible artifacts can be canonicalized and become reproducible after canonicalization.
@@ A.2. Implementation-Build-Java-Vendor
This attribute records the vendor of the JDK used to build the JAR file. There are 17 different open source distributions of JDK  [^39] https://sdkman.io/jdks/ - GraalVM, OpenJDK, Temurin to name a few. Although all of them follow the same Java specification by Oracle, the implementation details could differ.
@@ A.3. Build-Jdk or Implementation-Build-Java-Version
Example Artifact camel-cloud-3.13.0.jar of Maven release org.apache.camel:camel-cloud:3.13.0 is built with different versions of JDK. The reference version of the JAR file has the Build-Jdk attribute set to 1.8.0\_292 while the rebuild version of the JAR file has the Build-Jdk attribute set to 1.8.0\_422. The difference in the JDK version is due to the update number (Clark and Reinhold, 2014).
@@ A.11. Order of values for attributes
Solution The attributes where this problem exists are Include-Resource, Private-Package, and Provide-Capability. The solution to depend upon the latest version of maven-bundle-plugin as it makes the order deterministic by sorting the values  [^43] https://github.com/apache/felix-dev/commit/d885d99a6a16660f655a4fd18e8a1a39beef0a15.
@@ B.2. Timestamp
Solution The timestamp should either be set to the value of SOURCE\_DATE\_EPOCH or removed from the pom.properties file.
@@ B.3. Eclipse properties
If a Maven module is built using, m2e - Maven build tool for Eclipse IDE, it adds some properties to the pom.properties file other than the ones added by standalone Maven build tool.
@@ C.2. Addition or removal of components
Solution The build script here  [^46] https://github.com/pmd/pmd/blob/main/.ci/build.sh#L81 shows ad-hoc configuration of the build process. This should not be done as it hinders automation when checking for reproducibility. The fix here is to stick to conventional builds.
@@ C.3. Modification of components
C.6. System Package Data Exchange (SPDX)
SPDX is a standard to document the licenses of the components used to build the software. It has non-deterministic attributes like created and licenseListVersion which can cause unreproducibility. created attribute is used to indicate the time when the SPDX file is created. licenseListVersion attribute is used to indicate the version of the SPDX license list used to document the licenses of the components. This list is updated externally and can vary with time.
Example We only see one instance of unreproducibility that falls under this category. Maven release org.apache.commons:commons-parent:60 has an artifact commons-parent-60.spdx.json where value of created and licenseListVersion attribute is different across different builds.
Solution These attributes should be removed before comparison as they are non-deterministic.
@@ D.1. Files are removed or added
Example Maven release commons-daemon:commons-daemon:1.4.0 has an unreproducible artifact commons-daemon-1.4.0-bin-windows.zip which does not contain 3 Windows executable binary in the rebuild version of the archive. There is custom non-Maven command that has to be executed during the build process to generate and place these binaries under the build folder. Since the command is not documented in the buildspec file, the rebuild version of the archive does not contain these binaries.
@@ E.1. Debug information
This corresponds to the information in the JVM bytecode that is used for debugging the Java program. There are three types of debug information in the JVM bytecode - SourceFile, LineNumberTable, and LocalVariableTable.
@@ E.2. Optimizations or de-optimizations
  • ldc\_w instructions are replaced with ldc instructions. Example: org.codehaus.mojo:jaxb2-maven-plugin:3.1.0.
@@ E.3. Refactoring
Apart from order, we observe 2 types of refactoring that did not change the order.
@@ E.5. Build time generated/modified code
Example Maven release org.apache.helix:zookeeper-api:1.0.3 has an artifact zookeeper-api-1.0.3.jar. The lambda function changes from lambda\\$parseRoutingData\\$19 to lambda\\$parseRoutingData\\$1.
@@ F.3. Shell scripts
Solution The shell script should respect the Maven property \${project.build.outputTimestamp} or the standard SOURCE\_DATE\_EPOCH environment variable to set the value of the timestamp in the shell script.
@@ F.4. Binaries
Solution The build tool for binary should respect SOURCE\_DATE\_EPOCH environment variable to set the value of the timestamp in the binary.
@@ F.5. SBOM
Unreproducibility due to attribute timestamp in the SBOM as explained in section C.4.
@@ F.7. File timestamps
Files have timestamp for creation or modification time as shown in section D.4.
@@ F.8. MANIFEST
Attribute Bnd-LastModified tends to be non-deterministic as documented in section A.10.
@@ G.7. Attribute values in MANIFEST
Some values are comma separated lists whose order can differ as shown in section A.11.
@@ Appendix I Optimizations not canonicalized by jNorm
The third case that jNorm is unable to canonicalize the implementation of string concatenation when its operands require a cast to String. For example, "ErrordeterminingJDBCtypeforcolumn"+column+".Cause:"+e is a string concatenation expressions in artifact mybatis-3.5.16.jar. jNorm identifies that there are two ways to implement string concatenation (Shipilev, 2024) - using StringBuilder (JDK < 9) and using invokedynamic (JDK $\geq$ 9) instruction and canonicalizes the expression to the second one. However, it fails to canonicalize the string conversion of e which is an instance of java.sql.SQLException before concatenation using invokedynamic.
@@
jNorm is unable to handle differences in how outer class reference is handled in the bytecode. A synthetic field called this\\$0 is created in the bytecode to refer to the outer class from inner class. The reference version of com.google.guava:guava:32.0.1-jre is built using custom JDK which omits this synthetic field  [^73] https://bugs.openjdk.org/browse/JDK-8271717 as it is not needed. We also observe that patch versions of JDK 17 handle this differently. They refer to the outer class reference using either this.this\\$0 or this\\$0. The former adds another getfield instruction to the bytecode. psi-probe-tomcat85-3.7.0-tests.jar is an example of artifact that contains this difference.

### v2

@@ 8.3. Fixing Unreproducibility
Ren et al. (Ren et al., 2022, 2019, 2018) propose techniques to localize and fix unreproducibility in Debian packages. They localize the file in the Debian package that causes unreproducibility and then suggest a patch based on a history of patches for fixing unreproducible builds. This approach is specific for Debian packages and does not apply to Java artifacts. Moreover, their approach to detect unreproducibility is different from ours as mentioned in section˜8.2.
@@
Xiong et al. (Xiong et al., 2022) classifies fixing unreproducibility into three categories — 1) remediation by patching source code, third-party dependency, or build script 2) controlling which means interception of non-deterministic build instructions and returning pre-defined values for them 3) interpretation which means documenting the reason for unreproducibility. The authors perform these fixes automatically using a tool called JavaBEPFix which is not available to the public. Hence, we cannot evaluate the effectiveness of their tool on our dataset.
@@
Dietrich et al. (Dietrich et al., 2024b) show that only a 25% (out of 14,156) of the binaries in Maven ecosystem are actually bit by bit identical. These binaries come from four different registries — Maven, Google Assured Open Source Software, RedHat, and Oracle build-from-source. They propose four levels of binary equivalence to compare binaries which becomes increasingly lenient as we go from level 1 to level 4. First is to compare binaries bit by bit and this is the strictest level of binary equivalence check. Second is to compare binaries and ignore differences that have no semantic effect (for example, @Deprecated). OSS-Rebuild fits into the second level of checking binary equivalence. Third is to convert binaries into a common format and then compare them. jNorm fits into the third level of checking binary equivalence Finally, fourth is to compare binaries using a similarity metric. In the paper, they use tlsh (Oliver et al., 2013) to compare hash distance between two binaries; lesser the hash distance, more similar the binaries are. This work is similar to ours as they also evaluate effectiveness of jNorm over thousands of pairs of JVM bytecode (Dietrich et al., 2024a) to fix unreproducibility issues. However, we consider more causes of unreproducibility other than ones in JVM bytecode.
Finally, strip-nondeterminism  [^38] https://salsa.debian.org/reproducible-builds/strip-nondeterminism is a tool the Debian project that strips out timestamps and file system order from archives. However, it only takes care of these two causes of unreproducibility and does not cover the other causes of unreproducibility that we have discussed in this paper.
@@ 9. Conclusion
In this paper, we have presented the first large-scale study of build unreproducibility in Java. We have identified the six main causes of unreproducibility with underlying root causes: 1) build manifests, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps. Leveraging the same dataset, we have evaluated the effectiveness of canonicalization as a mitigation strategy to fix unreproducibility. Our results have shown that 26.89% of unreproducible artifacts can be canonicalized and become reproducible after canonicalization.
@@ A.2. Implementation-Build-Java-Vendor
This attribute records the vendor of the JDK used to build the JAR file. There are 17 different open source distributions of JDK  [^39] https://sdkman.io/jdks/ — GraalVM, OpenJDK, Temurin to name a few. Although all of them follow the same Java specification by Oracle, the implementation details could differ.
@@ A.3. Build-Jdk or Implementation-Build-Java-Version
Example Artifact camel-cloud-3.13.0.jar of Maven release org.apache.camel:camel-cloud:3.13.0 is built with different versions of JDK. The reference version of the JAR file has the Build-Jdk attribute set to 1.8.0_292 while the rebuild version of the JAR file has the Build-Jdk attribute set to 1.8.0_422. The difference in the JDK version is due to the update number (Clark and Reinhold, 2014).
@@ A.11. Order of values for attributes
Solution The attributes where this problem exists are Include-Resource, Private-Package, and Provide-Capability. The solution is to depend upon the latest version of maven-bundle-plugin as it makes the order deterministic by sorting the values  [^43] https://github.com/apache/felix-dev/commit/d885d99a6a16660f655a4fd18e8a1a39beef0a15.
@@ B.2. Timestamp
Solution The timestamp should either be set to the value of SOURCE_DATE_EPOCH or removed from the pom.properties file.
@@ B.3. Eclipse properties
If a Maven module is built using, m2e — Maven build tool for Eclipse IDE, it adds some properties to the pom.properties file other than the ones added by standalone Maven build tool.
@@ C.2. Addition or removal of components
Solution The build script here [^46] https://github.com/pmd/pmd/blob/7979570d39909195e173de05450fb1acf8b3eec5/.ci/build.sh#L81 shows ad-hoc configuration of the build process. This should not be done as it hinders automation when checking for reproducibility. The fix here is to stick to conventional builds.
@@ C.3. Modification of components
C.6. External Reference
This attribute indicates external systems which are related to the software. For example, version control system, issue tracking system, documentation, etc.
Example Gradle release org.jreleaser:jreleaser-mattermost-java-sdk:1.6.0 has an SBOM that only shows VCS reference for the components in the rebuilt version of the SBOM.
Solution The external reference should be removed via canonicalization.
C.7. Description of components
The description attribute is a free-form text that can be used to describe the component.
Example Gradle release org.jreleaser:jreleaser-nexus2-java-sdk:1.6.0 has an SBOM that only shows description for the components in the rebuilt version of the SBOM.
Solution The description of the components should be removed via canonicalization.
C.8. License
The difference in licence appears in Software Package Data Exchange (SPDX) format. SPDX is a standard to document the licenses of the components used to build the software. Apart from timestamps, it has a non-deterministic attribute licenseListVersion which can cause unreproducibility. licenseListVersion attribute is used to indicate the version of the SPDX license list used to document the licenses of the components. This list is updated externally and can vary with time.
Example Maven release org.apache.commons:commons-parent:60 has an artifact commons-parent-60.spdx.json where value of timestamp and licenseListVersion attribute is different across different builds.
Example Gradle release org.jreleaser:jreleaser-mattermost-java-sdk:1.6.0 has an SBOM that captures the license URLs for more components.
Solution License attribute should be removed before comparison as they are non-deterministic.
@@ D.1. Files are removed or added
Example Maven release commons-daemon:commons-daemon:1.4.0 has an unreproducible artifact commons-daemon-1.4.0-bin-windows.zip which does not contain three Windows executable binary in the rebuild version of the archive. There is custom non-Maven command that has to be executed during the build process to generate and place these binaries under the build folder. Since the command is not documented in the buildspec file, the rebuild version of the archive does not contain these binaries.
@@ E.1. Debug information
This corresponds to the information in the JVM bytecode that is used for debugging the Java program. There are three types of debug information in the JVM bytecode — SourceFile, LineNumberTable, and LocalVariableTable.
@@ E.2. Optimizations or de-optimizations
  • ldc_w instructions are replaced with ldc instructions. Example: org.codehaus.mojo:jaxb2-maven-plugin:3.1.0.
@@ E.3. Refactoring
Apart from order, we observe two types of refactoring that did not change the order.
@@ E.5. Build time generated/modified code
Example Maven release org.apache.helix:zookeeper-api:1.0.3 has an artifact zookeeper-api-1.0.3.jar. The lambda function changes from lambda\$parseRoutingData\$19 to lambda\$parseRoutingData\$1.
@@
Example Gradle release io.opentelemetry.instrumentation:opentelemetry-instrumentation-api:2.4.0 has an artifact opentelemetry-javaagent-2.4.0.jar.diffoscope.json which records MethodParameters attribute only the in reference version of the bytecode to allow reflective access to the name of the arguments.
@@ F.3. Shell scripts
Solution The shell script should respect the Maven property \${project.build.outputTimestamp} or the standard SOURCE_DATE_EPOCH environment variable to set the value of the timestamp in the shell script.
@@ F.4. Binaries
Solution The build tool for binary should respect SOURCE_DATE_EPOCH environment variable to set the value of the timestamp in the binary.
@@ F.5. SBOM
Unreproducibility due to attribute timestamp in the SBOM as explained in section˜C.4.
@@ F.7. File timestamps
Files have timestamp for creation or modification time as shown in section˜D.4.
@@ F.8. MANIFEST
Attribute Bnd-LastModified tends to be non-deterministic as documented in section˜A.10.
@@ G.7. Attribute values in MANIFEST
Some values are comma separated lists whose order can differ as shown in section˜A.11.
@@ Appendix I Optimizations not canonicalized by jNorm
The third case that jNorm is unable to canonicalize the implementation of string concatenation when its operands require a cast to String. For example, "ErrordeterminingJDBCtypeforcolumn"+column+".Cause:"+e is a string concatenation expressions in artifact mybatis-3.5.16.jar. jNorm identifies that there are two ways to implement string concatenation (Shipilev, 2024) — using StringBuilder (JDK < 9) and using invokedynamic (JDK $\geq$ 9) instruction and canonicalizes the expression to the second one. However, it fails to canonicalize the string conversion of e which is an instance of java.sql.SQLException before concatenation using invokedynamic.
@@
jNorm is unable to handle differences in how outer class reference is handled in the bytecode. A synthetic field called this\$0 is created in the bytecode to refer to the outer class from inner class. The reference version of com.google.guava:guava:32.0.1-jre is built using custom JDK which omits this synthetic field  [^73] https://bugs.openjdk.org/browse/JDK-8271717 as it is not needed. We also observe that patch versions of JDK 17 handle this differently. They refer to the outer class reference using either this.this\$0 or this\$0. The former adds another getfield instruction to the bytecode. psi-probe-tomcat85-3.7.0-tests.jar is an example of artifact that contains this difference.